# CMRNext: Camera to LiDAR Matching in the Wild for Localization and Extrinsic Calibration

### v3

@@ IV-J Extrinsic Calibration — Temporal Aggregation
where is the number of frames used for aggregation and is the quaternion corresponding to the -th frame. can be computed using the Singular Value Decomposition (SVD) of . For the median rotation, we evaluated the method proposed in [67], which starts from an initial guess, set to the average quaternion defined above, and iteratively updates the guess until convergence. However, due to the iterative nature of the method, the runtime was too high (the evaluation was not complete even after 24 hours) and therefore we did not include the median rotation in the results. Finally, the mode rotation does not require any special consideration, as the mode of a set of quaternions is simply the quaternion with the highest frequency in the set. For computing the mode of both translation and rotation, we first round all values to a given number of decimals, specifically two for translations and four for rotations.
@@
The multi-frames results reported in Table VIII show that the performance of all tendency measures decreases as the number of frames used for aggregation decreases, as hypothesized. Moreover, the mode consistently outperforms the mean measures on all experiments. The median aggregation achieves the best translation error on the left camera but decreases significantly on the right camera, where the mode achieves the best results. When using all frames of the testing sequence (4541 frames), the aggregated mode extrinsic calibration achieves a translation error of and a rotation error of on the left camera, and and on the right camera, setting the new state of the art for targetless LiDAR-camera extrinsic calibration.
@@ IV-K Extrinsic Calibration — Generalization
We evaluate the generalization ability of CMRNext on the extrinsic calibration task on our three in-house platforms detailed in Section IV-A: a self-driving perception car, a quadruped robot, and a UAV. To quantitatively evaluate our proposed method, we first calibrated the extrinsic parameters of each platform using a simple manual calibration procedure. In particular, we recorded a calibration sequence by moving a planar calibration board in front of the platform and then we manually selected the corners of the calibration board in both the camera images and the LiDAR scans. Finally, based on these correspondences, we used PnP with RANSAC followed by a Levenberg-Marquardt optimization based on the inliers of RANSAC. We compare the translation and rotation errors of CMRNext on the three platforms against the manual calibration. In this experiment, since the number of frames in each dataset is relatively low (around 20 frames), we use 30 different random initial noise poses for each frame and we report the median frame-wise errors and the error of the pose computed using temporal aggregation across all frames. The results reported in Table IX show that our method with temporal aggregation achieves rotation errors lower than on all three platforms and translation errors lower than on the car and UAV platforms. The translation error on the quadruped is considerably higher than those on the other platforms. However, this higher error might derive from an inaccurate manual calibration, as we have no way to guarantee the accuracy of our manual calibration procedure.
@@ IV-L Hyperparameter Optimization
The default hyperparameter values used for all the experiments reported in previous sections were optimized using the hyperparameter optimization method BOHB [68], which combines Bayesian optimization and Hyperband. In particular, we used the implementation provided with the Weigths&Biases platform [69]. The hyperparameters optimized were the learning rate, the batch size, the number of Fourier frequencies used for the positional encoding, the parameters of the occlusion filter, the optimizer, the learning rate scheduler, and the maximum depth at which a point in the LiDAR image is kept. Table X shows the search space used for each hyperparameter and the optimal value found by the hyperparameter optimization.
@@ IV-M Runtime Analysis
In this experiment, we compare the runtime of CMRNext with existing camera-agnostic methods CMRNet++ [26] and I2D-Loc [62]. Additionally, we compare the performance and runtime of different algorithms to solve the PnP problem, most of which implemented in the OpenCV library [70], including EPnP [47], SQPnP [71], our GPU-based implementation of EPnP + RANSAC. All runtimes were computed for the first iteration only on a workstation with 32GB DDR3 RAM, an Intel i7-4790K CPU, and an NVIDIA RTX 2080 Ti GPU. The results reported in Table XI show that CMRNext with the GPU-based PnP solver achieves the fastest runtime among all methods. Moreover, it is possible to further improve the performance of CMRNext by performing a Levenberg–Marquardt (LM) optimization on the inliers found by RANSAC, at the cost of increased runtime.
| Pose Estimator | Transl. | Rot. | Runime [s] | |
@@
(C) and (G) represent the CPU- and GPU-based implementation, respectively. LM represents an additional Levenberg–Marquardt optimization on the inliers found by RANSAC.
V Conclusions

### v4

@@ IV-J Extrinsic Calibration — Temporal Aggregation
where is the number of frames used for aggregation and is the quaternion corresponding to the -th frame. can be computed using the Singular Value Decomposition (SVD) of . For the median rotation, we evaluated the method proposed in [70], which starts from an initial guess, set to the average quaternion defined above, and iteratively updates the guess until convergence. However, due to the iterative nature of the method, the runtime was too high (the evaluation was not complete even after 24 hours), and therefore, we did not include the median rotation in the results. Finally, the mode rotation does not require any special consideration, as the mode of a set of quaternions is simply the quaternion with the highest frequency in the set. For computing the mode of both translation and rotation, we first round all values to a given number of decimals, specifically two for translations and four for rotations.
@@
The multi-frames results reported in Table VIII show that the performance of all tendency measures decreases as the number of frames used for aggregation decreases, as hypothesized. Moreover, the mode consistently outperforms the mean measures in all experiments. The median aggregation achieves the best translation error on the left camera but decreases significantly on the right camera, where the mode achieves the best results. When using all frames of the testing sequence (4541 frames), the aggregated mode extrinsic calibration achieves a translation error of and a rotation error of on the left camera, and and on the right camera, setting the new state of the art for targetless LiDAR-camera extrinsic calibration.
@@ IV-K Extrinsic Calibration — Generalization
We evaluate the generalization ability of CMRNext on the extrinsic calibration task on our three in-house platforms detailed in Section IV-A: a self-driving perception car, a quadruped robot, and a UAV. To quantitatively evaluate our proposed method, we first calibrated the extrinsic parameters of each platform using a simple manual calibration procedure. In particular, we recorded a calibration sequence by moving a planar calibration board in front of the platform and then we manually selected the corners of the calibration board in both the camera images and the LiDAR scans. Finally, based on these correspondences, we used PnP with RANSAC, followed by a Levenberg-Marquardt optimization based on the inliers of RANSAC. We compare the translation and rotation errors of CMRNext on the three platforms against the manual calibration. In this experiment, since the number of frames in each dataset is relatively low (around 20 frames), we use 30 different random initial noise poses for each frame, and we report the median frame-wise errors and the error of the pose computed using temporal aggregation across all frames. The results reported in Table IX show that our method with temporal aggregation achieves rotation errors lower than on all three platforms and translation errors lower than on the car and UAV platforms. The translation error on the quadruped is considerably higher than those on the other platforms. However, this higher error might derive from an inaccurate manual calibration, as we have no way to guarantee the accuracy of our manual calibration procedure.
@@ IV-L Hyperparameter Optimization
The default hyperparameter values used for all the experiments reported in previous sections were optimized using the hyperparameter optimization method BOHB [71], which combines Bayesian optimization and Hyperband. In particular, we used the implementation provided with the Weights&Biases platform [72]. The hyperparameters optimized were the learning rate, the batch size, the number of Fourier frequencies used for the positional encoding, the parameters of the occlusion filter, the optimizer, the learning rate scheduler, and the maximum depth at which a point in the LiDAR image is kept. Table X shows the search space used for each hyperparameter and the optimal value found by the hyperparameter optimization.
@@ IV-M Runtime Analysis
In this experiment, we compare the runtime of CMRNext with existing camera-agnostic methods CMRNet++ [27] and I2D-Loc [65]. Additionally, we compare the performance and runtime of different algorithms to solve the PnP problem, most of which are implemented in the OpenCV library [73], including EPnP [50], SQPnP [74], and our GPU-based implementation of EPnP + RANSAC. All runtimes were computed only for the first iteration on a workstation with 32GB DDR3 RAM, an Intel i7-4790K CPU, and an NVIDIA RTX 2080 Ti GPU. For the iterative refinement approach described in Section III-E, the runtime scales linearly with the number of iterations. The results reported in Table XI show that CMRNext with the GPU-based PnP solver achieves the fastest runtime among all methods. Moreover, it is possible to further improve the performance of CMRNext by performing a Levenberg–Marquardt (LM) optimization on the inliers found by RANSAC at the cost of increased runtime.
| Pose Estimator | Transl. | Rot. | Runtime [s] | |
@@
(C) and (G) represent the CPU- and GPU-based implementation, respectively. LM represents an additional Levenberg–Marquardt optimization on the inliers found by RANSAC. All runtimes are computed for a single iteration.
V Limitations
CMRNext has three primary limitations. First, because it is trained exclusively on autonomous driving datasets, its generalization ability is likely to be limited in other environments, such as off-road or mining environments. Second, while our method is faster than existing camera-agnostic approaches, it is not yet capable of real-time performance. Lastly, similar to most existing methods, the training paradigm requires ground truth poses as well as correct intrinsic camera parameters, which are typically provided in publicly available datasets, but add an additional constraint to the approach.
VI Conclusions